# Nauplius Optimisation for Autonomous Hydrodynamics

### v1

@@ V-B Drift and Update Equations
The drift term $\gamma U(x_{i,t})$ represents the’ core capability of NOAH for autonomous hydrodynamics, enabling agents to exploit ocean currents rather than resist them. Unlike traditional swarm algorithms that treat environmental forces as disturbances that must be overcome, NOAH agents actively use the hydrodynamic flow field $U(x)$ to improve exploration efficiency and reduce energy consumption. When $U(x_{i,t})$ points toward promising regions, agents accelerate their convergence; when currents flow away from optimal areas, agents can still navigate effectively through the combination of gradient following and colony attraction. This autonomous adaptation to hydrodynamic conditions mirrors how barnacle nauplii exploit favourable flow structures for settlement [37], transforming environmental constraints into navigational advantages. The parameter $\gamma\in[0.3,0.8]$ controls the degree of current exploitation, allowing NOAH to operate effectively in diverse marine environments, from gentle coastal flows to strong oceanic currents.
@@ V-C Settlement Decision and Irreversibility
where $\Delta f_{i}=\tilde{f}_{N}-f(x_{i,t})$ represents the local fitness advantage, $d_{i}=\min_{k\in C_{t}}\|x_{i,t}-c_{k}\|$ is the distance to the nearest colony, $\kappa(x_{i,t})=\|\nabla U(x_{i,t})\|_{2}$ measures flow shear (turbulence), $\rho_{i}$ indicates local crowding and $E_{i}\in[0,1]$ encodes energy reserves. The parameters $\lambda_{1},\dots,\lambda_{5}$ control the relative importance of each factor, while $d_{0}$ and $\kappa_{0}$ provide normalisation.
@@ V-D Colony Dynamics and Communication
Over time, colony strength $S_{k}$ is dynamically adjusted based on reinforcement from local successes and penalties from overcrowding, subject to the probability of communication success. Let $\overline{\text{rank}(-f)}_{\text{near }c_{k}}$ denote the average rank of f-f among agents near colony $k$, and let $\overline{\rho}_{\text{near }c_{k}}$ be the average agent density in its vicinity. The update rule incorporates communication probability as
@@
where $\mu\in(0,1]$ is a learning rate and $\nu$ scales the crowding penalty. The probability of communication success $p_{k}^{\text{comm}}\in[0,1]$ is calculated as the average of $p_{k}(x)$ on all nearby agents, representing the reliability of communication between the colony $k$ and its surrounding agents. When $p_{k}^{\text{comm}}=1$, the colony receives full updates; when $p_{k}^{\text{comm}}=0$, the colony strength decays without reinforcement, simulating communication failure. This update increases colony strength if nearby agents continue to produce high-quality solutions and the area is not over-saturated. In contrast, if local performance declines or density grows excessively, strength decreases. Colonies thus behave as living entities: reinforced by successful discoveries and weakened by redundancy. This mechanism parallels the reinforcement of pheromones and evaporation in the optimisation of the ant colony [12].
@@ V-E Complexity and Convergence Considerations
The computational complexity introduces additional overhead compared to standard swarm optimisers such as PSO [10] and ACO [12]. This is due to the colony field computation and settlement decision processes. Let $N$ denote the number of agents and $C$ the number of active colonies at time $t$.
Each iteration requires $O(N)$ operations for agent state updates, $O(N\cdot C)$ for colony field computation $G_{\Phi}(x)$ at each agent location, $O(N)$ for settlement probability evaluation using the logistic model [71], and $O(C)$ for colony strength updates based on pheromone reinforcement mechanisms [12]. In the worst case where $C=O(N)$ (all agents eventually settle), the total complexity per iteration is $O(N^{2})$.
@@
The communication architecture is specifically designed to minimise bandwidth requirements in underwater acoustic networks, where communication is expensive and unreliable [2]. Unlike approaches where all agents continuously exchange information, the algorithm leverages colonies to reduce communication load through a hierarchical structure. Free agents broadcast locally (positions, fitness values) and listen for colony beacons, while anchored agents act as cluster heads, broadcasting colony strength and location information. During iteration, the complexity of communication is $O(N+C)$ messages, with $O(N)$ local broadcasts from free agents and $O(C)$ colony broadcasts. This represents a significant reduction compared to fully-connected swarm algorithms and enables multi-hop acoustic networking through anchored agents.
@@
The algorithm implements a progressive freeze strategy. It begins with an initial exploration phase where all agents explore the search space using velocity updates that incorporate inertia [61], random exploration with Gaussian noise [72] and gradient estimation [73]. During the discovery phase, high-quality regions are identified and secured through anchoring. The exploitation phase guides the remaining free agents to promising areas through colony fields based on attraction-repulsion dynamics [54] [55]. The system converges to a metastable state where all agents are anchored, creating a distributed sensor network.
@@ V-G1 Hydrodynamic Drift Adaptation
11:   $\tilde{v}_{i,t+1}=\omega v_{i,t}+\eta\xi_{i,t}+\beta(-\hat{\nabla}f(x_{i,t}))+\gamma U(x_{i,t})+\delta G_{\Phi}(x_{i,t})$
@@ VI-A Anchoring Accuracy Validation
Results demonstrate 86% success (86/100 runs). Distance from optimum ranged 0.021-0.485 units (mean: 0.194±0.104), with fitness error 0.0004-0.235 (mean: 0.049±0.046). Best performance achieved position (0.020,0.008)(-0.020,0.008) with fitness 0.9996-0.9996, demonstrating near-perfect accuracy. Colony formation averaged 5.11 per run, validating reliability for permanent deployment.
VI-B Convergence Behavior Analysis
Figure 7: NOAH convergence behavior across four benchmark functions over 200 iterations. Solid lines represent mean best fitness, shaded regions indicate standard deviation across 30 independent runs. The progressive improvement demonstrates effective exploration-exploitation balance through irreversible anchoring.
@@ IX Future Research Directions
Based on the successful validation of the’ anchoring capabilities of NOAH, several promising research directions emerge that could significantly advance underwater swarm robotics. These directions are categorised into immediate next steps and long-term research opportunities.
@@ X Conclusion
The integration of biological inspiration with algorithmic design represents a novel approach to marine robotics. By translating barnacle nauplii behaviours into a unified mathematical framework, NOAH bridges the gap between terrestrial swarm algorithms and marine-specific requirements, providing a new paradigm for underwater swarm coordination. The validated foundation demonstrates the’ potential of NOAH for marine exploration, conservation, and research applications, offering significant advancement in underwater robotics through domain-specific algorithmic solutions.
@@ Acknowledgements
The authors extend their gratitude to Dr. Tony de Souza-Daw, Dr. Fernando Galetto, and Adam Console for their continued support and encouragement that helped shape the direction and quality of this research. The contributions provided were instrumental in strengthening both the technical and practical aspects of this study.
@@ Appendix A Mathematical Notation
$\text{rank}(x)$ Function that ranks values from best (1) to worst

### v2

@@ V-B Drift and Update Equations
The drift term $\gamma U(x_{i,t})$ represents the core capability of NOAH for autonomous hydrodynamics, enabling agents to exploit ocean currents rather than resist them. Unlike traditional swarm algorithms that treat environmental forces as disturbances that must be overcome, NOAH agents actively use the hydrodynamic flow field $U(x)$ to improve exploration efficiency and reduce energy consumption. When $U(x_{i,t})$ points toward promising regions, agents accelerate their convergence; when currents flow away from optimal areas, agents can still navigate effectively through the combination of gradient following and colony attraction. This autonomous adaptation to hydrodynamic conditions mirrors how barnacle nauplii exploit favourable flow structures for settlement [larsson2016instantaneous], transforming environmental constraints into navigational advantages. The parameter $\gamma\in[0.3,0.8]$ controls the degree of current exploitation, allowing NOAH to operate effectively in diverse marine environments, from gentle coastal flows to strong oceanic currents.
@@ V-C Settlement Decision and Irreversibility
where $\Delta f_{i}=\tilde{f}_{N}-f(x_{i,t})$ represents the local fitness advantage, $d_{i}=\min_{k\in C_{t}}\|x_{i,t}-c_{k}\|$ is the distance to the nearest colony, $\kappa(x_{i,t})=\|\nabla U(x_{i,t})\|_{2}$ measures flow shear (turbulence), $\rho_{i}$ indicates local crowding and $E_{i}\in[0,1]$ encodes energy reserves. The parameters $\lambda_{1},\dots,\lambda_{5}$ control the relative importance of each factor, while $d_{0}$ and $\kappa_{0}$ are normalising constants. Here, $\text{rank}(\Delta f_{i})$ denotes the scalar rank of agent $i$ among all agents (1 for best fitness, 2 for second-best, etc.), not the matrix rank.
@@ V-D Colony Dynamics and Communication
Over time, colony strength $S_{k}$ is dynamically adjusted based on reinforcement from local successes and penalties from overcrowding, subject to the probability of communication success. Let $\overline{\text{rank}(-f)}_{\text{near }c_{k}}$ denote the average of these scalar ranks for agents near colony $k$, and let $\overline{\rho}_{\text{near }c_{k}}$ be the average agent density in its vicinity. The update rule incorporates communication probability as
@@
where $\mu\in(0,1]$ is a learning rate and $\nu$ scales the crowding penalty. The probability of communication success $p_{k}^{\text{comm}}\in[0,1]$ is calculated as the average of $p_{k}(x)$ on all nearby agents, representing the reliability of communication between the colony $k$ and its surrounding agents. When $p_{k}^{\text{comm}}=1$, the colony receives full updates; when $p_{k}^{\text{comm}}=0$, the colony strength decays without reinforcement, simulating communication failure. This update increases colony strength if nearby agents continue to produce high-quality solutions and the area is not over-saturated. In contrast, if local performance declines or density grows excessively, strength decreases. Colonies thus behave as living entities: reinforced by successful discoveries and weakened by redundancy. This mechanism parallels the reinforcement of pheromones and evaporation in the optimisation of the ant colony [dorigo2006ant].
@@ V-E Complexity and Convergence Considerations
The computational complexity introduces additional overhead compared to standard swarm optimisers such as PSO [kennedy1995particle] and ACO [dorigo2006ant]. This is due to the colony field computation and settlement decision processes. Let $N$ denote the number of agents and $C$ the number of active colonies at time $t$.
Each iteration requires $O(N)$ operations for agent state updates, $O(N\cdot C)$ for colony field computation $G_{\Phi}(x)$ at each agent location, $O(N)$ for settlement probability evaluation using the logistic model [fagerland2012generalized], and $O(C)$ for colony strength updates based on pheromone reinforcement mechanisms [dorigo2006ant]. In the worst case where $C=O(N)$ (all agents eventually settle), the total complexity per iteration is $O(N^{2})$.
@@
The communication architecture is specifically designed to minimise bandwidth requirements in underwater acoustic networks, where communication is expensive and unreliable [connor2021current]. Unlike approaches where all agents continuously exchange information, the algorithm leverages colonies to reduce communication load through a hierarchical structure. Free agents broadcast locally (positions, fitness values) and listen for colony beacons, while anchored agents act as cluster heads, broadcasting colony strength and location information. During iteration, the complexity of communication is $O(N+C)$ messages, with $O(N)$ local broadcasts from free agents and $O(C)$ colony broadcasts. This represents a significant reduction compared to fully-connected swarm algorithms and enables multi-hop acoustic networking through anchored agents.
@@
The algorithm implements a progressive freeze strategy. It begins with an initial exploration phase where all agents explore the search space using velocity updates that incorporate inertia [shi1998modified], random exploration with Gaussian noise [wang2003stochastic] and gradient estimation [zhang2012local]. During the discovery phase, high-quality regions are identified and secured through anchoring. The exploitation phase guides the remaining free agents to promising areas through colony fields based on attraction-repulsion dynamics [gazi2004class] [gazi2006coordination]. The system converges to a metastable state where all agents are anchored, creating a distributed sensor network.
@@ V-G1 Hydrodynamic Drift Adaptation
11:   $\tilde{v}_{i,t+1}=\omega v_{i,t}+\eta\xi_{i,t}+\beta(-g(x_{i,t}))+\gamma U(x_{i,t})+\delta G_{\Phi}(x_{i,t})$
@@ VI-A Anchoring Accuracy Validation
Results demonstrate 86% success (86/100 runs). Distance from optimum ranged from 0.021 to 0.485 units (mean: 0.194±0.104), with fitness error ranging from 0.0004 to 0.235 (mean: 0.049±0.046). Best performance achieved position (0.020,0.008)(-0.020,0.008) with fitness 0.9996-0.9996, demonstrating near-perfect accuracy. Colony formation averaged 5.11 per run, validating reliability for permanent deployment.
VI-B Convergence Behaviour Analysis
Figure 7: NOAH convergence behaviour across four benchmark functions over 200 iterations. Solid lines represent mean best fitness, shaded regions indicate standard deviation across 30 independent runs. The progressive improvement demonstrates effective exploration-exploitation balance through irreversible anchoring.
@@ IX Future Research Directions
Based on the successful validation of the anchoring capabilities of NOAH, several promising research directions emerge that could significantly advance underwater swarm robotics. These directions are categorised into immediate next steps and long-term research opportunities.
@@ X Conclusion
The integration of biological inspiration with algorithmic design represents a novel approach to marine robotics. By translating barnacle nauplii behaviours into a unified mathematical framework, NOAH bridges the gap between terrestrial swarm algorithms and marine-specific requirements, providing a new paradigm for underwater swarm coordination. The validated foundation demonstrates the potential of NOAH for marine exploration, conservation, and research applications, offering significant advancement in underwater robotics through domain-specific algorithmic solutions.
@@ Acknowledgements
The authors extend their gratitude to Dr. Guang Deng, Dr. Tony de Souza-Daw, Dr. Fernando Galetto, and Adam Console for their continued support and encouragement that helped shape the direction and quality of this research. The contributions provided were instrumental in strengthening both the technical and practical aspects of this study.
@@ Appendix A Mathematical Notation
$\text{rank}(x)$ Scalar rank of a value within a set (1 = best), not matrix rank
$p_{k}^{\text{comm}}$ Average communication-success probability for colony $k$
$\overline{\text{rank}(-f)}_{\text{near }c_{k}}$ Average scalar rank of f-f for agents near colony $k$